What freezing point depression would have resulted
Answer #1

m = 0.59 g of ehxane

disolved in 20.6 g of PDB

Kf = -7.10

so..

kg of solvent (PDB) = 20.6/1000 = 0.0206 kg

Mw of hexane = 86.18 g/mol

mol = mass/MW = 0.59/86.18 = 0.0068461

molality = mol/kg solvent = 0.0068461/0.0206 = 0.33233

so

dTF = -K*m = -7.10*0.33233

dTF = -2.3595 °C lower than normal freezing point